# Adaptive Multilevel Splitting: First Application to Rare Event Derivative Pricing

### v2

@@ 4.4 Analysis of option pricing results
Performance is evaluated in terms of computational time (horizontal axis) and relative accuracy (vertical axis), defined as
$\frac{\sqrt{\mathrm{Var}}}{\mathrm{Mean}},$
with the mean and variance estimated over $50$ independent runs. Figures 3 and 4 summarise these results: each curve represents one simulation method for a given contract, and each point on a curve corresponds to a different computational budget. Times are shown on a logarithmic scale, so vertical separations between curves can be interpreted directly as orders of magnitude in speed-up. The underlying numerical values used to generate these plots are reported in Tables 69.
@@
We denote by MCA Monte Carlo with antithetic variates, and by MCV Monte Carlo with the optimised lower-bound control variate of Fusai and Kyriakou [FusaiKyriakou2016].
@@
In Figure 4, an additional benchmark is included: the control variate method derived from the optimised lower bound introduced by [FusaiKyriakou2016]. This provides a further point of comparison, confirming that even when specialised variance–reduction techniques tailored to the payoff structure are incorporated, AMS remains the most computationally efficient method across the accuracy levels examined.
@@
4.4.1 Extreme case analysis
@@ 4.4.2 Multi-asset digital option analysis
ensures that the sufficient condition of [brehier2016unbiasedness, Theorem 3.2] holds and the AMS estimator remains unbiased.
@@ 5.2 Future developments
A first direction is risk management, where AMS could improve the computation of tail risk measures. Value-at-Risk (VaR), defined as the loss threshold exceeded with small probability, is a rare event problem. Existing Monte Carlo and importance sampling approaches are widely used [hong2014monte, sun2010asymptotic]; AMS offers the potential for more accurate and efficient estimates, relevant for stress testing and regulatory capital.
@@
A third avenue is methodological. Rough volatility models such as Bergomi [gatheral2017rough] pose challenges because fractional Brownian motion violates the Markov property central to AMS. One possible solution is a lifted Markovian approximation embedding the non-Markovian dynamics in a higher-dimensional state space [zhu2021markovian], potentially extending AMS to this class of models.
@@ Appendix A C++ implementation with R interface via Rcpp
Prior to this work, no R package provided AMS functionality for financial applications. To fill this gap, a dedicated implementation was developed in C++ [press2007numerical] with an R interface via Rcpp [eddelbuettel2013seamless].
@@
The Rcpp interface exposes all algorithmic parameters and diagnostics within the R environment, while computationally intensive tasks remain in C++. This design combines the usability of R with near native performance, enabling practical deployment of AMS in quantitative finance. The complete implementation is distributed as the R package amsSim [mio], available on CRAN.

### v3

@@ 4.4 Analysis of option pricing results
Performance is evaluated in terms of computational time (horizontal axis) and relative accuracy (vertical axis), defined as $\frac{\sqrt{\mathrm{Var}}}{\mathrm{Mean}}$, with the mean and variance estimated over $50$ independent runs. Figures 3 and 4 summarise these results: each curve represents one simulation method for a given contract, and each point on a curve corresponds to a different computational budget. Times are shown on a logarithmic scale, so vertical separations between curves can be interpreted directly as orders of magnitude in speed-up. The underlying numerical values used to generate these plots are reported in Tables 69.
@@
We denote by MCA Monte Carlo with antithetic variates, and by MCV Monte Carlo with the optimised lower-bound control variate of Fusai and Kyriakou [19].
@@
In Figure 4, an additional benchmark is included: the control variate method derived from the optimised lower bound introduced by [19]. This provides a further point of comparison, confirming that even when specialised variance–reduction techniques tailored to the payoff structure are incorporated, AMS remains the most computationally efficient method across the accuracy levels examined.
@@
4.4.1 Theoretical analysis of Greek estimation in adaptive frameworks
Estimating sensitivities (Greeks) within the AMS framework presents specific challenges related to the adaptive nature of the algorithm. This section contrasts the Finite Difference approach with direct differentiation methods, highlighting the structural bias affecting the latter.
Finite Difference benchmark. The Finite Difference (FD) method serves as a robust baseline. For a parameter $\theta$ and perturbation $h$, the estimator is
$\partial_{\theta}V\approx\frac{\hat{V}(\theta+h)-\hat{V}(\theta)}{h}$
Although FD typically suffers from variance inflation as $h\to 0$, the variance reduction achieved by AMS on the price estimator $\hat{V}$ mitigates this instability. This makes FD a viable strategy without requiring prohibitive computational budgets.
Bias in pathwise differentiation. The conditional pathwise method [27] was implemented to address the non-differentiability of binary payoffs. This technique smoothens the discontinuity by conditioning on the state of the process just prior to maturity (at $T-\Delta t$) and integrating out the final Brownian increment. This effectively replaces the indicator $\mathbf{1}_{\{S_{T}>K\}}$ with the conditional probability of exercise, a smooth function that permits the interchange of differentiation and expectation.
While this approach yields substantial variance reduction, it exhibits a systematic negative bias in the AMS setting. The bias stems from the dependence of the AMS weight on model parameters. Considering the estimator $\hat{V}_{\text{AMS}}=W_{\text{AMS}}\times\hat{Y}_{N}$, the total derivative requires the product rule:
$\frac{\partial\hat{V}_{\text{AMS}}}{\partial\theta}=W_{\text{AMS}}\frac{\partial\hat{Y}_{N}}{\partial\theta}+\hat{Y}_{N}\frac{\partial W_{\text{AMS}}}{\partial\theta}.$ (7)
Standard pathwise schemes operate under a “frozen weight” approximation, computing only the first term. However, the stopping time $Q$ in AMS is parameter-dependent: for a Call option, an increase in $S_{0}$ reduces the expected iterations ($\partial Q/\partial S_{0}<0$), implying $\partial W_{\text{AMS}}/\partial S_{0}>0$. Neglecting this term causes the observed underestimation.
Computational trade-off. This issue mirrors the path degeneracy problem documented in Sequential Monte Carlo (SMC) literature [34]. Jasra and Del Moral demonstrated that the resampling mechanism, central to both SMC and AMS, introduces discrete updates that disrupt the smooth dependence of trajectories on parameters. Consequently, standard gradient estimators fail to account for the changes in particle genealogy induced by parameter perturbations.
Retrieving unbiased gradients via the Likelihood Ratio Method (LRM) necessitates integrating over all possible ancestral lineages to capture the weight sensitivity, imposing a computational complexity of $\mathcal{O}(N^{2})$. For the large population sizes required in rare event simulation ($N\sim 10^{5}$), such quadratic cost is prohibitive. Conversely, linear complexity approximations exist but suffer from asymptotic bias.
Finite Differences represent the most effective trade-off: they avoid the systematic bias of pathwise smoothing and the prohibitive cost of LRM. While subject to discretization error, they implicitly capture the sensitivity of the stopping time $Q$ while maintaining linear $\mathcal{O}(N)$ complexity.
4.4.2 Numerical stress test
@@ 4.4.3 Multi-asset digital option analysis
ensures that the sufficient condition of [8, Theorem 3.2] holds and the AMS estimator remains unbiased.
@@ 5.2 Future developments
A first direction is risk management, where AMS could improve the computation of tail risk measures. Value-at-Risk (VaR), defined as the loss threshold exceeded with small probability, is a rare event problem. Existing Monte Carlo and importance sampling approaches are widely used [32, 50]; AMS offers the potential for more accurate and efficient estimates, relevant for stress testing and regulatory capital.
@@
A third avenue is methodological. Rough volatility models such as Bergomi [22] pose challenges because fractional Brownian motion violates the Markov property central to AMS. One possible solution is a lifted Markovian approximation embedding the non-Markovian dynamics in a higher-dimensional state space [54], potentially extending AMS to this class of models.
@@ Appendix A C++ implementation with R interface via Rcpp
Prior to this work, no R package provided AMS functionality for financial applications. To fill this gap, a dedicated implementation was developed in C++ [44] with an R interface via Rcpp [18].
@@
The Rcpp interface exposes all algorithmic parameters and diagnostics within the R environment, while computationally intensive tasks remain in C++. This design combines the usability of R with near native performance, enabling practical deployment of AMS in quantitative finance. The complete implementation is distributed as the R package amsSim [28], available on CRAN.